Exercise 6-11 Future and present value [LO6-3, 6-6, 6-7] Answer each of the following independent questions Alex Meir recently won a lottery and has the option of receiving one of the following three prizes: (1) $82,000 cash immediately, (2) $30,000 cash immediately and a six-period annuity of $9,000 beginning one year from today, or (3) a six-period annuity of $17,000 beginning one year from today. (FV of $1, PV of $1, FVA of $1, PVA of $1. FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) I. Assuming an interest rate of 7%, determine the present value for the above options, which option should Alex choose? 2. The Weimer Corporation wants to accumulate a sum of money to repay certain debts due on December 31, 2027. Weimer will make annual deposits of $160,000 into a special bank account at the end of each of 10 years beginning December 31, 2018, Assuming that the bank account pays 8% interest compounded annually, what will be the fund balance after the last payment is made on December 31, 2027? Complete this question by entering your answers in the tabs below Required 1 Required 2 Assuming an interest rate of 7%, determine the present value for the above options, which option should Alex choose? (Round your final answers to nearest whole dollar amount.) Annuity Payment Immediate Cash PV Annuity PV Option Option 1 Option 2 Option 3 Which option should Alex choose? < Required! Required 2>

media%2F77c%2F77c9ed3e-32c8-4a41-b39e-5f
Answer #1

1.

Annuity payment PV annuity Immediate cash PV option
Option 1 0.00 0.00 + 82,000.00 = 82,000
Option 2 9,000.00 42,898.86 + 30,000.00 = 72,899
Option 3 17,000.00 81,031.18 + 0.00 = 81,031
Which option should Alex chose The first option

Calculations and explanations: Here we will use  present value of annuity factor i.e. PVA $1. From the table we get present value of annuity for 6 years and 7% = 4.76654.

Thus for option 2 present value of annuity of $9000 = 9000*4.76654 = 42,898.86 and for option 3 pv of annuity = 17000*4.76654 = 81031.18

2.

Table or calculator function FVA
Payment 160,000.00
n= 10.00
i= 8.00
Future value 2,317,856

Explanation: FVA for 8% and 10 years = 14.4866. Thus future value = 160000*14.4866 = 2317856